OpenStudy (mellamokatie):

8. (2 × 10^3)(4 × 10^–5) A. 8 × 10^–2 B. 0.8 × 10^–2 C. 8 × 10^–1 D. 8 × 10^2

OpenStudy (rcp031):

2*4*10^(3-5) gives answer

OpenStudy (will.h):

Is that x or a multiply sign

OpenStudy (mellamokatie):

Wait let me see

OpenStudy (mellamokatie):

Yes it is

OpenStudy (shadowlegendx):

\[(2 \times 10^3)(4\times 10^{-5})\]

OpenStudy (mellamokatie):

Yes perfect thank you

OpenStudy (mellamokatie):

I think I know what you would do can I say?

OpenStudy (mellamokatie):

Do you multiply the exponents 3(-5)?

OpenStudy (will.h):

Okay so this case require to simplify the expression 2 * 4 = 8 And 10^3 * 10^-5 here we can simply that by adding the exponent because in multiply property we add the exponents for the expressions with the same base So 10^3 * 10^-5 = 10^ 3+ (-5) So 10^-2 So the final expression would be 8 x 10^-2

OpenStudy (rcp031):

no, actually it is like \[1000/100000=1/100\]

OpenStudy (will.h):

Let me know if you have additional questions

OpenStudy (mellamokatie):

Thank you